### v1

# KAN KAN Buff Signed Graph Neural Networks?

Abstract
Graph Representation Learning aims to create embeddings for nodes and edges, capturing their features and interconnections. Graph Neural Networks (GNNs) have excelled in this task, leveraging neural networks to model complex graph relationships. Recently, the Kolmogorov-Arnold Neural Network (KAN) emerged as an alternative to Multi-Layer Perceptron (MLP), showing improved accuracy and interpretability with fewer parameters. While KANs have been integrated into unsigned GNNs, their application in signed GNNs remains unexplored. This paper integrates KAN into Signed Graph Convolutional Networks (SGCNs) to evaluate its performance on signed graphs where edges have positive or negative signs. We empirically assess KAN-enhanced SGCNs (KASGCN) on downstream tasks such as signed community detection and link sign prediction to enhance the embedding quality in signed networks. Considering the variability in the results indicated by the relatively large standard deviation, KASGCN demonstrates competitive performance with, or similar to, the vanilla SGCN in the evaluated downstream tasks, and its effectiveness is context-dependent (signed graph and parameters…etc.).
Kolmogorov-Arnold Neural Networks, Signed Network, Graph Representation Learning, Graph Neural Networks
1 Introduction
Graph Representation Learning is the process of fitting embeddings of nodes/edges based on the features and connections of the graph. These embeddings capture the local and global features of the nodes and edges in the original graph and are used in many downstream tasks such as node classification and link prediction [Xu, 2021]. One of the most prominent methods for learning graph embeddings is Graph Neural Networks (GNNs). They have shown remarkable effectiveness in capturing graphs’ complex relationships and structures [Fan et al., 2019, Bessadok et al., 2023]. GNNs utilize neural networks to gather and process information from a node’s surrounding nodes, gradually creating more sophisticated embeddings through successive layers.
Recently, the Kolmogorov-Arnold Neural Network (KAN) emerged as an alternative to Multi-Layer Perceptron (MLP) [Liu et al., 2024] and has demonstrated promising results in various benchmarks. Inspired by the Kolmogorov-Arnold representation theorem, KANs have learnable activation functions on edges, replacing fixed activation functions on nodes. This method has improved accuracy and interpretability, often surpassing MLPs with fewer parameters and more compact network architectures. Consequently, there have been multiple attempts to explore, validate, and integrate KANs in a variety of applications and architectures, which encompasses graph neural networks [De Carlo et al., 2024, Kiamari et al., 2024, Bresson et al., 2024] in a bid to achieve better performance and improve the quality of the embeddings. However, all the related work focuses on employing KANs in unsigned GNNs, and there is no research done regarding the effect of KANs on signed GNNs, which are designed to learn embeddings for signed graphs where edges can have positive or negative signs indicating different types of interactions.
In this paper, we adopt the most basic variant of signed GNNs proposed by Derr et al. [Derr et al., 2018], which is a Signed Graph Convolutional Neural Network (SGCN) for simplicity and as a baseline. This variant of GNN was the first to adapt the Graph Convolutional Neural Network for signed networks based on the balance theory. It is a pivotal concept that explains the evolution of attitudes within networks. Established by Heider [Abelson and Rosenberg, 1958] and subsequently formalized mathematically by Harary, who introduced $k$-way balancing [Cartwright and Harary, 1956, Harary and Cartwright, 1968], balance theory has applications in predicting edge sentiment, content and product recommendations, and anomaly detection in multiple domains [Derr et al., 2020, Garimella et al., 2021, Interian et al., 2022, Amelkin and Singh, 2019]. A signed network is considered strongly balanced if every fundamental cycle is positive (the number of negative edges for every cycle is even). We are the first to integrate KAN (with its variants) into signed graph neural networks and empirically evaluate its impact on performance on downstream tasks (signed clustering and link sign prediction) for signed networks. In addition, this study aims to explore the potential of KANs in increasing the expressive power of the embeddings generated by signed graph neural networks.
The paper is organized as follows: Section 2 discusses the related work regarding KAN integration into several architectures and their uses. In Section 3, we present the methodology of our work. In Section 4, we empirically evaluate the KAN-integrated signed GNN and its variants in several tasks, including efficiency examination and embedding comparison. We structure this section by presenting the research question, hypothesis, results, and observations. In Section 5, we summarize our findings.
2 Related Work
Cheon et al. [Cheon, 2024] introduced a pioneering method that integrates Kolmogorov-Arnold Networks (KANs) with various pre-trained Convolutional Neural Network models, such as VGG16 and MobileNetV2, for Remo, the sensing scene classification using the Eurostat dataset. This approach achieved high accuracy with fewer epochs and parameters than traditional Multi-Layer Perceptrons without KAN integration. Bresson et al. proposed to integrate the KAN layer into unsigned graph neural networks (GNNs) and evaluated its performance in the node and graph classification tasks [Bresson et al., 2024]. Li et al. utilized Fourier series-based KAN’s power to optimize GNNs, and the evaluation took place on datasets for molecular property prediction [Li et al., 2024]. KANs were also synergized with science by using a pykan package that contains many functionalities, like compiling symbolic formulas into KANs. Based on these tools, KANs’ ability to detect many physical and constitutive laws is demonstrated [Liu et al., 2024]. In addition, Dong et al. evaluated the performance of KANs on time series data and found that KANs can match or even surpass the performance of MLP across 128 different time series datasets [Dong et al., 2024]. KANs have been explored for data representation using autoencoders and compared against traditional Convolutional Neural Networks (CNNs) on datasets such as MNIST, SVHN, and CIFAR-10. And it is shown that KAN-based autoencoders deliver competitive reconstruction accuracy [Moradi et al., 2024]. Lu et al. assessed using KANs for fraud detection and concluded that their performance varies depending on the context [Lu and Zhan, 2024]. Finally, the study in [Anonymous, 2024] presented a proof of concept for employing KANs in graphs to predict molecules’ binding affinity to protein targets. The model in this study did not attain state-of-the-art performance, but it emphasized its promising method in computational drug discovery.
3 Methodology
3.1 Kolmogorov-Arnold Networks Background
According to the Kolmogorov-Arnold representation theorem [Liu et al., 2024], any multivariate function can be formulated as a combination of continuous univariate functions. It can be represented as:
$f(x_{1},x_{2},\ldots,x_{n})=\sum_{i=1}^{2d+1}\alpha_{i}\left(\sum_{j=1}^{d}\phi_{ij}(x_{j})\right)$ (1)
where $\alpha$ and $\phi$ are the univariate functions and $d$ is the dimension of the input. A univariate and non-linear spline function is applied for each input dimension. The output of each is aggregated. The main hyperparameters for the spline function are the spline order (degree of B-splines) and the grid size (number of intervals to approximate the real function). In the original paper [Liu et al., 2024], KANs demonstrated greater expressiveness than standard multi-layer perceptrons (MLPs), and KANs surpassed MLPs with significantly fewer parameters. In MLPs, weight adjustments occur globally, which can lead to forgetting previously learned data when new data is introduced. Conversely, in KANs, modifying spline control points impacts only local regions (locality property), thereby preserving previous training. $\phi(x)$ is formulated as:
$\phi(x)=w_{b}b(x)+w_{s}\ \text{spline}(x).$ (2)
$b(x)=\text{silu}(x)=\frac{x}{1+e}$ (3)
$\text{spline}(x)=\sum_{i}c_{i}B_{i}(x)$ (4)
where $b(x)$ is the SiLU activation function, $\text{spline}(x)$ is a linear combination of B-splines, $c_{i}$ is a trainable parameter and $w_{b}$ and $w_{s}$ are trainable and control the magnitude of the activation function. $w_{s}=1$ and $w_{b}$ is initialized using Xavier initialization.
3.2 Adaptation of Signed Graph Convolutional Network to Kolmogorov-Arnold Signed Graph Convolutional Network
Derr. et al. [Derr et al., 2018] employ balance theory to aggregate features from a node’s neighbors appropriately. Each node in the signed network has a dual representation (positive and negative representation). During message passing, to update the positive representation of a node, the positive embeddings are aggregated from the positive neighbors, and the negative embeddings are aggregated from the negative neighbors to be concatenated with each other along its positive representation. On the other hand, for updating the negative representation of a node, the negative embeddings are aggregated from the positive neighbors, and the positive embeddings are aggregated from the negative neighbors to be concatenated with each other along its negative representation. Finally, the positive and negative embeddings are concatenated to obtain the final embedding. The signed GCN formulation is as follows:
For $l=1:$
$h_{i}^{B(1)}=\sigma(W^{B(1)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (5)
$h_{i}^{U(1)}=\sigma(W^{U(1)}[\sum_{k\in N_{i}^{+}}{\frac{h_{k}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (6)
For $l>1:$
$h_{i}^{B(l)}=\sigma(W^{B(l)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{B(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{U(l-1)}}{|N_{i}^{-}|}},h_{i}^{B(l-1)}])$ (7)
$h_{i}^{U(l)}=\sigma(W^{B(l)}[\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{U(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{B(l-1)}}{|N_{i}^{-}|}},h_{i}^{U(l-1)}])$ (8)
where $l$ is the layer number, $h_{i}^{B(l)}$ and $h_{i}^{U(l)}$ are the positive and negative representations of node $i$ in layer $l$ respectively, $\sigma$ is an activation function, $W^{B(l)}$ and $W^{U(l)}$ are the weight matrices for updating the positive and negative representations respectively, $N_{i}^{+}$ and $N_{i}^{-}$ are the sets of positive and negative neighbors of $i$ respectively. Inspired by [Bresson et al., 2024], we integrate the Kolmogorov-Arnold layer (KAN) in the signed graph convolutional neural network (SGCN) by replacing the activation function $\sigma$ and the weight matrix $W$ with the KAN layer in the formulation. This improvement is easily integrated as the KAN layers have trainable parameters and introduce non-linearity, superseding the roles of the activation function and the weight matrix. Therefore, the Kolmogorov-Arnold signed graph convolutional network (KASGCN) can be formulated as follows:
For $l=1:$
$h_{i}^{B(1)}=\phi^{B(1)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (9)
$h_{i}^{U(1)}=\phi^{U(1)}([\sum_{k\in N_{i}^{+}}{\frac{h_{k}^{0}}{|N_{i}^{+}|}},h_{i}^{0}])$ (10)
For $l>1:$
$h_{i}^{B(l)}=\phi^{B(l)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{B(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{U(l-1)}}{|N_{i}^{-}|}},h_{i}^{B(l-1)}])$ (11)
$h_{i}^{U(l)}=\phi^{U(l)}([\sum_{j\in N_{i}^{+}}{\frac{h_{j}^{U(l-1)}}{|N_{i}^{+}|}},\sum_{k\in N_{i}^{-}}{\frac{h_{k}^{B(l-1)}}{|N_{i}^{-}|}},h_{i}^{U(l-1)}])$ (12)
where $\phi^{B(l)}$ and $\phi^{U(l)}$ here represent the KAN layer used to update the positive and negative representations, respectively. This layer can be any state-of-the-art variant such as OriginalKAN $\phi_{o}$ [Liu et al., 2024], FourierKAN $\phi_{F}$ [Xu et al., 2024], LaplaceKAN $\phi_{l}$ [Yin et al., 2024], and WaveletKAN $\phi_{w}$ [Bozorgasl and Chen, 2024]. We later experimented with these variants to determine how each performed when integrated into SGCN.
4 Empirical Evaluation
4.1 Setup
The operating system used for the experiments is Linux Ubuntu 20.04.3, running on the 11th Gen Intel(R) Core(TM) i9-11900K @ 3.50GHz with 16 physical cores. It is one socket, with two threads per core and eight cores per socket. The architecture is X86_x64. The GPU is Nvidia GeForce RTX 3070 and has 8GB of memory. Its driver version is 495.29.05, and the CUDA version is 11.5. The cache configuration is L1d : 384 KiB, L1i : 256 KiB, L2 : 4 MiB, L3 : 16 MiB. The CPU op is 32-bit and 64-bit.
4.2 Parameters
Each KAN layer consists of a grid of splines. The degree of each spline is the spline order, and the number of splines for each function is known as grid size. For the implementation of B-splines and other variants of KAN, we rely on the publicly available implementations in [Blealtan, 2024] and [Yin et al., 2024], respectively. The parameters used in the experiments are: layers = [32,32], weight_decay = $10^{-5}$, learning_rate = 0.001, lamb = 1.0, epochs = 1000, seed = 42, reduction_iterations = 10, reduction_dimensions = 15, spectral_features = True, norm = True, norm_embed = True, grid_size = 5, spline_order = 3, scale_noise = 0.1, scale_base = 1.0, scale_spline = 1.0, base_activation = torch.nn.SiLU, grid_eps = 0.02, grid_range= [-1, 1], Kmeans++ is used for clustering and with default parameters. Each experiment is repeated 10 times and averaged. The test size split for each positive and negative for link sign prediction (using multinomial logistic regression) is 0.2. t-SNE parameters are random_state = 0, n_iter = 1000, metric = ’cosine’. Preprocessing of the signed network includes Removing duplicates and keeping a first edge, treating neutral edges as positive, dropping self-loops, reindex nodes from 0 to $n-1$ where $n$ is the number of nodes in the signed network. A tanh activation function is applied after passing the embedding to the KAN layer.
4.3 Evaluation Metrics
We use two metrics to gauge the quality of our clustering assignments: fraction of positive edges within clusters and fraction of negative edges between clusters). They are the following:
$pos_{in}=\frac{pos_{within}}{pos_{between}+pos_{within}}$ (13)
$neg_{out}=\frac{neg_{between}}{neg_{within}+neg_{between}}$ (14)
where $pos_{within}$ and $pos_{between}$ are the number of positive edges within and between clusters, respectively. $neg_{within}$ and $neg_{between}$ are the negative edges within and between clusters. We rely on the AUC score that summarizes the GNN’s effectiveness across all possible classification thresholds to evaluate the link sign prediction performance. In addition, we use the F1 score for this as well, which is as follows:
[…]
where $Precision$ is the ratio of true positive instances to the sum of true positive and false positive instances. And $Recall$ is the ratio of true positive instances to the sum of true positive and false negative instances. We use seconds to measure and compare the training time of the graph neural network (GNN) models. The average cosine similarity is employed to measure the similarity between the embeddings generated by two different GNNs for each node in the signed network, and the formula is:
Avg Cosine $\text{Avg Cosine Sim}=\frac{1}{N}\sum_{i\in N}\frac{\mathbf{e}_{i}^{A}\cdot\mathbf{e}_{i}^{B}}{\|\mathbf{e}_{i}^{A}\|\|\mathbf{e}_{i}^{B}\|}$ (16)
where $N$ is the set of all nodes in the network, $\mathbf{e}_{i}^{A}$ and $\mathbf{e}_{i}^{B}$ are the embeddings of node $i$ generated from GNN $A$ and GNN $B$ respectively.
4.4 Signed Graphs
Table 1 describes the Konect and other signed graphs and their characteristics [Kunegis, 2013]. BitcoinAlpha is a user-user trust/distrust network from the Bitcoin Alpha platform for trading bitcoins. BitcoinOTC is a user-user trust/distrust network from the Bitcoin OTC platform for trading Bitcoins. WikiRFA describes voting information for electing Wikipedia managers [He et al., 2022]. WikiElec is the network of users from the English Wikipedia that voted for and against each other in admin elections. In the Chess network, each vertex is a chess player, and a directed edge represents a game with the white player having an outgoing edge and the black player having an ingoing edge. The weight of the edge represents the outcome. Congress is a signed network where vertices are politicians speaking in the United States Congress, and a directed edge denotes that a speaker mentions another speaker. PPI models the protein-protein interaction network [He et al., 2022]. Preprocessing removes duplicates (keep first), treats neutral edges as positive, removes self-loops, and reindexes nodes from 0 to the maximum number of nodes in the signed graph.
Table 1: Konect plus WikiRFA and PPI properties. Vertices, edges, and cycles represent the signed graph, not just the largest connected component (LCC). The rest of the metrics correspond to the LCC.
Graph Vertices Edges Cycles Density Triads Avg Deg. Median deg. Max deg. % of $e^{-}$
BitcoinAlpha 3,775 14,120 10,346 $<$ 0.01 22,153 7.48 2 511 8.39
BitcoinOTC 5,875 21,489 15,615 0.01 33,493 7.31 2 795 13.57
WikiRFA 7,634 167,936 160,303 $<$ 0.01 1,240,033 43.99 13 1,223 22.98
WikiElec 7,066 100,667 93,602 $<$ 0.01 607,279 28.49 4 1,065 21.94
Chess 7,115 55,779 48,665 $<$ 0.01 108,584 15.67 7 181 24.15
Congress 219 521 303 0.021 212 4.71 3 33 20.34
PPI 3,058 11,860 8,803 $<$ 0.01 3,837 3.87 2 55 32.5
4.5 Tasks
4.5.1 Community Detection
  • Research Question: Does incorporating the KAN layer into the signed graph convolutional network improve the clustering quality when using Kmeans++ on the embeddings?
  • Hypothesis: KASGCN learns representations that enhance the clustering of vertices by minimizing the negative edges between different clusters and maximizing the number of positive edges within the same cluster.
  • Results: The clustering results are summarized in Table 2, Table 3, and Table 4 for $K=5$, $K=10$, $K=15$ respectively where $K$ is the predefined number of clusters parameter as input for the Kmeans++ algorithm.
    Table 2: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=5$).
    GNN SGCN KASGCN
    Quality $pos$ $neg$ $pos$ $neg$ Gain
    BitcoinAlpha 0.388 $\pm$ 0.0003 0.8015 $\pm$ 0.003 0.4093 $\pm$ 0.011 0.659 $\pm$ 0.05 -10.18%
    BitcoinOTC 0.513 $\pm$ 0.011 0.820 $\pm$ 0.054 0.523 $\pm$ 0.00033 0.707 $\pm$ 0.001 -7.70%
    WikiElec 0.507 $\pm$ 0.0002 0.773 $\pm$ 0.0004 0.5221 $\pm$ 0.004 0.784 $\pm$ 0.0009 2.03%
    WikiRFA 0.410 $\pm$ 0.0025 0.843 $\pm$ 0.0021 0.391 $\pm$ 0.0010 0.795 $\pm$ 0.0006 -5.34%
    Chess 0.434 $\pm$ 0.002 0.641 $\pm$ 0.014 0.409 $\pm$ 0.015 0.619 $\pm$ 0.034 -4.27%
    Congress 0.616 $\pm$ 0.028 0.961 $\pm$ 0.009 0.566 $\pm$ 0.029 0.9584 $\pm$ 0.017 -3.33%
    PPI 0.495 $\pm$ 0.001 0.709 $\pm$ 0.010 0.385 $\pm$ 0.084 0.87 $\pm$ 0.0109 4.23%
    Table 3: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=10$).
    GNN SGCN KASGCN
    Quality $pos$ $neg$ $pos$ $neg$ Gain
    BitcoinAlpha 0.275 $\pm$ 0.023 0.911 $\pm$ 0.0040 0.299 $\pm$ 0.006 0.877 $\pm$ 0.012 -0.84%
    BitcoinOTC 0.378 $\pm$ 0.020 0.929 $\pm$ 0.0045 0.421 $\pm$ 0.005 0.864 $\pm$ 0.051 -1.60%
    WikiElec 0.3260 $\pm$ 0.005 0.8956 $\pm$ 0.003 0.283 $\pm$ 0.020 0.909 $\pm$ 0.0022 -2.40%
    WikiRFA 0.2637 $\pm$ 0.004 0.910 $\pm$ 0.0013 0.270 $\pm$ 0.033 0.917 $\pm$ 0.04 1.13%
    Chess 0.254 $\pm$ 0.026 0.798 $\pm$ 0.038 0.22 $\pm$ 0.0054 0.888 $\pm$ 0.0022 5.30%
    Congress 0.497 $\pm$ 0.023 0.996 $\pm$ 0.0048 0.472 $\pm$ 0.050 0.992 $\pm$ 0.0074 -1.80%
    PPI 0.422 $\pm$ 0.0366 0.784 $\pm$ 0.011 0.364 $\pm$ 0.037 0.936 $\pm$ 0.0093 7.79%
    Table 4: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for SGCN and KASGCN node embeddings using Kmeans++ clustering ($K=15$).
    GNN SGCN KASGCN
    Quality $pos$ $neg$ $pos$ $neg$ Gain
    BitcoinAlpha 0.2266 $\pm$ 0.010 0.943 $\pm$ 0.0070 0.245 $\pm$ 0.0074 0.921 $\pm$ 0.006 -0.30%
    BitcoinOTC 0.335 $\pm$ 0.015 0.956 $\pm$ 0.005 0.380 $\pm$ 0.17 0.93 $\pm$ 0.014 1.47%
    WikiElec 0.249 $\pm$ 0.067 0.935 $\pm$ 0.002 0.227 $\pm$ 0.0090 0.942 $\pm$ 0.04 -1.20%
    WikiRFA 0.197 $\pm$ 0.0072 0.9407 $\pm$ 0.037 0.203 $\pm$ 0.0051 0.9470 $\pm$ 0.0050 1.08%
    Chess 0.195 $\pm$ 0.003 0.853 $\pm$ 0.0033 0.181 $\pm$ 0.0024 0.922 $\pm$ 0.0020 5.24%
    Congress 0.438 $\pm$ 0.027 0.997 $\pm$ 0.0045 0.392 $\pm$ 0.034 0.996 $\pm$ 0.0048 -3.27%
    PPI 0.333 $\pm$ 0.01 0.833 $\pm$ 0.0064 0.309 $\pm$ 0.031 0.959 $\pm$ 0.0049 8.74
  • Observations: Improvement in the clustering quality highly depends on the choice of K. For relatively higher values of K, Kmeans++ using embeddings from KASGCN seems to do slightly better. In comparison, for relatively lower values of K, Kmeans++ using embeddings from SGCN seems to do slightly better. Overall, the difference in the clustering quality between SGCN and KASGCN is negligible or meager except in PPI, where KASGCN outperforms SGCN regardless of K.
4.5.2 Link Sign Prediction
  • Research Question: Do the embeddings generated from integrating the KAN layer into SGCN increase the predictive power of logistic regression for sign classification?
  • Hypothesis: The expressiveness of KASGCN’s embeddings allows logistic regression to better predict the sign of the edge than that of SGCN.
  • Results: The link sign prediction results are presented in Table 5, where a Multinomial Logistic Regression is trained on the embeddings of both SGCN and KASGCN with random 0.2 test size splits over 10 runs.
    Table 5: Average AUC and F1 score over 10 runs for SGCN and KASGCN node embeddings using Multinomial Logistic Regression with random 0.2 test size split.
    GNN SGCN KASGCN Gain
    Metric AUC F1 AUC F1 AUC F1
    BitcoinAlpha 0.783 $\pm$ 0.014 0.721 $\pm$ 0.042 0.799 $\pm$ 0.022 0.690 $\pm$ 0.094 2.04% -4.29%
    BitcoinOTC 0.840 $\pm$ 0.005 0.792 $\pm$ 0.014 0.859 $\pm$ 0.0075 0.761 $\pm$ 0.068 2.26% -3.91%
    WikiElec 0.809 $\pm$ 0.0044 0.781 $\pm$ 0.0087 0.837 $\pm$ 0.0045 0.787 $\pm$ 0.032 3.46% 0.76%
    WikiRFA 0.820 $\pm$ 0.0031 0.7623 $\pm$ 0.014 0.830 $\pm$ 0.0037 0.759 $\pm$ 0.017 1.21% -0.434
    Chess 0.542 $\pm$ 0.005 0.612 $\pm$ 0.017 0.548 $\pm$ 0.0064 0.625 $\pm$ 0.041 1.10% 2.12%
    Congress 0.570 $\pm$ 0.055 0.63 $\pm$ 0.058 0.499 $\pm$ 0.046 0.602 $\pm$ 0.077 -12.29% -4.44%
    PPI 0.679 $\pm$ 0.015 0.614 $\pm$ 0.067 0.692 $\pm$ 0.015 0.654 $\pm$ 0.037 1.91% 6.51%
  • Observations: Logistic Regression with embeddings from KASGCN shows consistent improvement in AUC across most datasets, indicating better overall performance in distinguishing between classes. The highest AUC improvement is observed in the WikiElec dataset, with a 3.46% increase. The F1 score improvements are mixed. While there are gains in datasets like WikiElec (0.76%) and Chess (2.12%), there are notable declines in datasets like BitcoinAlpha (-4.29%) and BitcoinOTC (-3.91%). While Logistic Regression with embeddings from KASGCN generally enhances AUC, the impact on F1 scores is graph-dependent.
4.6 Efficiency Examination
  • Research Question: How efficient is KASGCN compared to SGCN?
  • Hypothesis: Incorporating the KAN layer into the Signed Graph Neural Network increases the model’s complexity, leading to longer execution times for training embeddings.
  • Results: We visualize the efficiency of training SGCN and KASGCN embeddings in Figure 1 with varying numbers of aggregator layers.
    Figure 1: The embedding generation time of SGCN and KASGCN with varying aggregator layers.
  • Observations: As expected, due to the spline function, KAGCN takes a significantly longer time to train and generate embeddings. Increasing the aggregation layers impacts the execution time for KASGCN but not SGCN.
4.7 Embeddings Comparison
  • Research Question: Given that the observed improvements are relatively modest, how similar do SGCN and KASGCN produce the embeddings? Are these slight enhancements possible due to the similarity in the embeddings generated by the two models?
  • Hypothesis: The embeddings generated by SGCN and KASGCN are similar.
  • Results: The embeddings for SGCN (Red) and KASGCN (Blue) are visualized using t-SNE for BitcoinAlpha, BitcoinOTC, and WikiElec in Figure 2. The average cosine similarity between the embeddings of SGCN and KASGCN for these three signed graphs is computed to be -0.02, 0.01, and 0.072, respectively.
    Figure 2: t-SNE visualizations of the embeddings generated by SGCN (Red) and KASGCN (Blue) for different signed graphs.
  • Observations: Similarities are near zero, which indicates orthogonality. The t-SNE scatter plots show that the red and blue points are significantly dissimilar, highlighting that the embeddings do not match across the board.
4.8 KAN Variants Trial
  • Research Question: Does integrating other KAN variants into SGCN perform just as well as integrating the original KAN layer?
  • Hypothesis: The embeddings generated by other SGCN integrated with other KAN variants outperform those generated by SGCN integrated with the original KAN layer in the community detection and link sign prediction tasks in the signed network.
  • Results: The comparison of the community detection results between all the variants of the Kolmogorov-Arnold Layer is presented in Table 6, Table 7, and Table 8 for $K=5$, $K=10$, $K=15$ respectively. Table 9 shows the link sign prediction results of the Kolmogorov-Arnold Layer variants where a Multinomial Logistic Regression is trained on the embeddings of both SGCN and KASGCN with random 0.2 test size splits over 10 runs.
    Table 6: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=5$).
    GNN KASGCN FourierKASGCN LaplaceKASGCN WaveletKASGCN
    Quality pos neg pos neg pos neg pos neg
    BitcoinAlpha 0.4093 $\pm$ 0.011 0.659 $\pm$ 0.05 0.37 $\pm$ 0.0003 0.766 $\pm$ 0.0007 0.462 $\pm$ 0.014 0.716 $\pm$ 0.019 0.3879 $\pm$ 0.0016 0.786 $\pm$ 0.0011
    BitcoinOTC 0.523 $\pm$ 0.00033 0.707 $\pm$ 0.001 0.515 $\pm$ 0.004 0.841 $\pm$ 0.0004 0.467 $\pm$ 0.019 0.721 $\pm$ 0.062 0.511 $\pm$ 0.0003 0.833 $\pm$ 0.0001
    WikiElec 0.5221 $\pm$ 0.004 0.784 $\pm$ 0.0009 0.546 $\pm$ 0.0021 0.7844 $\pm$ 0.0014 0.418 $\pm$ 0.0071 0.725 $\pm$ 0.003 0.472 $\pm$ 0.0 0.781 $\pm$ 0.0
    WikiRFA 0.391 $\pm$ 0.0010 0.795 $\pm$ 0.0006 0.425 $\pm$ 0.0003 0.807 $\pm$ 0.0003 0.409 $\pm$ 0.0062 0.708 $\pm$ 0.010 0.422 $\pm$ 0.019 0.853 $\pm$ 0.0016
    Chess 0.409 $\pm$ 0.015 0.619 $\pm$ 0.034 0.380 $\pm$ 0.091 0.759 $\pm$ 0.017 0.551 $\pm$ 0.007 0.451 $\pm$ 0.0078 0.345 $\pm$ 0.00043 0.717 $\pm$ 0.00039
    Congress 0.566 $\pm$ 0.029 0.9584 $\pm$ 0.017 0.554 $\pm$ 0.031 0.988 $\pm$ 0.0097 0.290 $\pm$ 0.054 0.647 $\pm$ 0.033 0.648 $\pm$ 0.057 0.977 $\pm$ 0.014
    PPI 0.385 $\pm$ 0.084 0.87 $\pm$ 0.0109 0.445 $\pm$ 0.015 0.781 $\pm$ 0.0027 0.838 $\pm$ 0.061 0.223 $\pm$ 0.101 0.365 $\pm$ 0.0023 0.810 $\pm$ 0.0067
    Table 7: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=10$).
    GNN KASGCN FourierKASGCN LaplaceKASGCN WaveletKASGCN
    Quality pos neg pos neg pos neg pos neg
    BitcoinAlpha 0.299 $\pm$ 0.006 0.877 $\pm$ 0.012 0.253 $\pm$ 0.003 0.896 $\pm$ 0.0031 0.352 $\pm$ 0.019 0.770 $\pm$ 0.0087 0.281 $\pm$ 0.0077 0.876 $\pm$ 0.0052
    BitcoinOTC 0.421 $\pm$ 0.005 0.864 $\pm$ 0.051 0.401 $\pm$ 0.017 0.922 $\pm$ 0.0086 0.370 $\pm$ 0.014 0.834 $\pm$ 0.015 0.4100 $\pm$ 0.0029 0.9246 $\pm$ 0.005
    WikiElec 0.283 $\pm$ 0.020 0.909 $\pm$ 0.0022 0.2741 $\pm$ 0.001 0.920 $\pm$ 0.00085 0.267 $\pm$ 0.013 0.823 $\pm$ 0.0027 0.252 $\pm$ 0.0036 0.910 $\pm$ 0.0018
    WikiRFA 0.270 $\pm$ 0.033 0.917 $\pm$ 0.04 0.250 $\pm$ 0.0033 0.930 $\pm$ 0.0012 0.400 $\pm$ 0.017 0.715 $\pm$ 0.031 0.269 $\pm$ 0.0016 0.924 $\pm$ 0.0076
    Chess 0.22 $\pm$ 0.0054 0.888 $\pm$ 0.0022 0.243 $\pm$ 0.0021 0.873 $\pm$ 0.0014 0.419 $\pm$ 0.013 0.592 $\pm$ 0.012 0.225 $\pm$ 0.0048 0.864 $\pm$ 0.0068
    Congress 0.472 $\pm$ 0.050 0.992 $\pm$ 0.0074 0.482 $\pm$ 0.037 1.0 $\pm$ 0.0 0.182 $\pm$ 0.016 0.845 $\pm$ 0.036 0.521 $\pm$ 0.037 0.995 $\pm$ 0.0066
    PPI 0.364 $\pm$ 0.037 0.936 $\pm$ 0.0093 0.353 $\pm$ 0.0254 0.884 $\pm$ 0.0059 0.732 $\pm$ 0.022 0.387 $\pm$ 0.035 0.289 $\pm$ 0.0087 0.922 $\pm$ 0.0083
    Table 8: Average $pos\_{in}$ and $neg\_{out}$ over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Kmeans++ clustering ($K=15$).
    GNN KASGCN FourierKASGCN LaplaceKASGCN WaveletKASGCN
    Quality pos neg pos neg pos neg pos neg
    BitcoinAlpha 0.245 $\pm$ 0.0074 0.921 $\pm$ 0.006 0.219 $\pm$ 0.0095 0.932 $\pm$ 0.0074 0.277 $\pm$ 0.024 0.837 $\pm$ 0.020 0.227 $\pm$ 0.0089 0.933 $\pm$ 0.0053
    BitcoinOTC 0.380 $\pm$ 0.17 0.93 $\pm$ 0.014 0.317 $\pm$ 0.012 0.965 $\pm$ 0.0041 0.316 $\pm$ 0.020 0.874 $\pm$ 0.011 0.337 $\pm$ 0.0136 0.954 $\pm$ 0.0041
    WikiElec 0.227 $\pm$ 0.0090 0.942 $\pm$ 0.04 0.214 $\pm$ 0.0081 0.946 $\pm$ 0.0042 0.189 $\pm$ 0.018 0.866 $\pm$ 0.010 0.166 $\pm$ 0.0081 0.937 $\pm$ 0.0012
    WikiRFA 0.203 $\pm$ 0.0051 0.9470 $\pm$ 0.0050 0.194 $\pm$ 0.0074 0.954 $\pm$ 0.0018 0.301 $\pm$ 0.0071 0.813 $\pm$ 0.0072 0.198 $\pm$ 0.0090 0.948 $\pm$ 0.0028
    Chess 0.181 $\pm$ 0.0024 0.922 $\pm$ 0.0020 0.191 $\pm$ 0.0035 0.919 $\pm$ 0.0020 0.374 $\pm$ 0.018 0.631 $\pm$ 0.019 0.178 $\pm$ 0.0048 0.912 $\pm$ 0.0026
    Congress 0.392 $\pm$ 0.034 0.996 $\pm$ 0.0048 0.4286 $\pm$ 0.029 0.999 $\pm$ 0.0029 0.149 $\pm$ 0.023 0.906 $\pm$ 0.029 0.439 $\pm$ 0.0300 1.0 $\pm$ 0.0
    PPI 0.309 $\pm$ 0.031 0.959 $\pm$ 0.0049 0.307 $\pm$ 0.023 0.909 $\pm$ 0.006 0.684 $\pm$ 0.026 0.458 $\pm$ 0.0396 0.251 $\pm$ 0.014 0.941 $\pm$ 0.0020
    Table 9: Average AUC and F1 score over 10 runs for KASGCN, FourierKASGCN, LaplaceKASGCN, and WaveletKASGCN node embeddings using Multinomial Logistic Regression with random 0.2 test size split.
    GNN KASGCN FourierKASGCN LaplaceKASGCN WaveletKASGCN
    Quality AUC F1 AUC F1 AUC F1 AUC F1
    BitcoinAlpha 0.799 $\pm$ 0.022 0.690 $\pm$ 0.094 0.796 $\pm$ 0.018 0.656 $\pm$ 0.19 0.582 $\pm$ 0.039 0.095 $\pm$ 0.302 0.782 $\pm$ 0.020 0.705 $\pm$ 0.090
    BitcoinOTC 0.859 $\pm$ 0.0075 0.761 $\pm$ 0.068 0.848 $\pm$ 0.0082 0.785 $\pm$ 0.082 0.513 $\pm$ 0.046 0.0020 $\pm$ 0.0044 0.855 $\pm$ 0.0088 0.773 $\pm$ 0.065
    WikiElec 0.837 $\pm$ 0.0045 0.787 $\pm$ 0.032 0.819 $\pm$ 0.014 0.793 $\pm$ 0.046 0.488 $\pm$ 0.024 0.00030 $\pm$ 0.0009 0.828 $\pm$ 0.015 0.800 $\pm$ 0.043
    WikiRFA 0.830 $\pm$ 0.0037 0.759 $\pm$ 0.017 0.830 $\pm$ 0.0040 0.782 $\pm$ 0.030 0.496 $\pm$ 0.02 0.0 $\pm$ 0.0 0.836 $\pm$ 0.0046 0.765 $\pm$ 0.0338
    Chess 0.548 $\pm$ 0.0064 0.625 $\pm$ 0.041 0.550 $\pm$ 0.0059 0.635 $\pm$ 0.068 0.508 $\pm$ 0.012 0.018 $\pm$ 0.035 0.552 $\pm$ 0.0049 0.610 $\pm$ 0.047
    Congress 0.499 $\pm$ 0.046 0.602 $\pm$ 0.077 0.544 $\pm$ 0.059 0.656 $\pm$ 0.074 0.493 $\pm$ 0.086 0.013 $\pm$ 0.042 0.563 $\pm$ 0.083 0.590 $\pm$ 0.069
    PPI 0.692 $\pm$ 0.015 0.654 $\pm$ 0.037 0.689 $\pm$ 0.019 0.661 $\pm$ 0.071 0.513 $\pm$ 0.030 0.0 $\pm$ 0.0 0.694 $\pm$ 0.016 0.641 $\pm$ 0.047
  • Observations: In the community detection task, identifying the best KASGCN variant is challenging, as performance appears to be influenced by both the K value and the signed graph. However, LaplaceKASGCN consistently underperforms across various graphs. For the link sign prediction task and across all datasets, LaplaceKASGCN also shows poor performance, which may not be suitable for this task. In contrast, FourierKASGCN and WaveletKASGCN demonstrate superior performance on most graphs, as evidenced by their higher F1 and AUC metrics.
5 Summary
In this study, we integrated the Kolmogorov-Arnold Neural Network into the Signed Graph Convolutional Network to investigate its effectiveness in several tasks, including community detection and link sign prediction in signed networks. Overall, our empirical evaluations demonstrated that KAN-enhanced SGCNs achieve comparable performance compared to vanilla SGCNs, considering the variability in the results indicated by the relatively large standard deviation, and it is context-dependent (signed graph, parameters..etc.). Our study provides a foundation and paves the way for further research concerning the applicability of KANs in various domains.